OpenStudy (anonymous):

Simplify. n 6 · n 5 ÷ n 4 · n 3 ÷ n 2 · n

OpenStudy (anonymous):

I assume the numbers are exponents? Such as: \[n^6*n^5÷n^4*n^3÷n^2*n\] Utilize the Order of Operations when solving. Remember these rules: When multiplying numbers of the same base, add exponents: \[n^1*n^2=n^3\] When dividing numbers of the same base, subtract the exponents: \[n^2/n^1=n^1\]